# CPT: Competence-progressive Training Strategy for Few-shot Node Classification

### v3

@@ 1 Introduction
In this paper, we propose Competence-Progressive Training strategy for few-shot node classification, named as CPT. CPT aims to provide tasks of different difficulty levels for meta-learners at different stages to adapt to their current abilities and alleviate the issue of suboptimal points. Specifically, CPT uses the two-stage strategy Bengio et al. (2009); Hu et al. (2022) to train a meta-learner. In the first stage, CPT follows previous research by randomly sampling nodes to construct meta tasks and then employs the default training strategy(episodic learning), i.e., simply minimizing the given supervised loss, to produce a strong base meta-learner to start with. Although the base meta-learner may perform well on easy tasks and has gained some knowledge, it no longer benefits or improves its abilities from these tasks, and it might converge to suboptimal points. So the second stage focuses on generating more difficult tasks and fine-tuning it. Specifically, CPT increases the difficulty of each task by dropping the edges of its graph. The underlying motivation for this approach is to create a larger number of ’tail nodes’ within the meta-training tasks. This alteration is intended to enrich the meta-learning model with additional ’tail node’ meta-knowledge, as these nodes, with their fewer connections, typically present unique challenges in learning processes. Finally, CPT finetunes the base meta-learner by minimizing the loss over the increased hard task data. CPT can be easily implemented in the existing meta-learner training pipeline on graphs, as demonstrated in Algorithm 1. Additionally, CPT is compatible with any GNN-based meta-learner and can be employed with any desired supervised loss function.
@@ 2 PRELIMINARIES
2.1 Problem Statement
Formally, let $G=(\mathcal{V},\mathcal{E},\mathcal{X})$ denote an attributed graph, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}$ is the set of edges, and $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the feature matrix of nodes with $d$ denoting the feature dimension. Moreover, we denote the entire set of node classes as $C$, which can be further divided into two categories: $C_{b}$ and $C_{n}$, where $C=C_{b}\cup C_{n}$ and $C_{b}\cap C_{n}=\emptyset$. Here $C_{b}$ and $C_{n}$ denote the sets of base and novel classes, respectively. It is worth mentioning that the number of labeled nodes in $C_{b}$ is sufficient, while it is typically small in $C_{n}$. Then we can formulate the studied problem of few-shot node classification as follows:
Few-shot Node Classification: Given an attributed graph $G=(\mathcal{V},\mathcal{E},\mathcal{X})$ our goal is to develop a machine learning model such that after training on labeled nodes in $C_{b}$ , the model can accurately predict labels for the nodes (i.e., query set $Q$) in $C_{n}$ with only a limited number of labeled nodes (i.e., support set $S$).
More specifically, if the support set $S$ contains exactly $K$ nodes for each of $N$ classes from $C_{n}$, and the query set $Q$ are sampled from these $N$ classes, the problem is called $N$-way $K$-shot node classification. Essentially, the objective of few-shot node classification is to learn a classifier that can be fast adapted to $C_{n}$ with only limited labeled nodes. Thus, the crucial part is to learn transferable knowledge from $C_{b}$ and generalize it to $C_{n}$.
2.2 Episodic Learning
@@ 4.3 Implementation Details
Our methodology is formulated utilizing the PyTorch [^1] https://pytorch.org/ framework, and all subsequent evaluations are executed on a Linux-based server equipped with 5 NVIDIA Tesla V100 GPUs. For the specific implementation setting, we set the number of training epochs $T$ as 2000, the weight decay rate is set as 0.0005, the loss weight $\gamma$ is set as 1, the graph layers are set as 2 and the hidden size is set as in $\left\{16,32\right\}$. The READOUT function is implemented as mean-pooling. The other optimal hyper-parameters were determined via grid search on the validation set: the learning rate was searched in $\left\{0.03,0.01,0.005\right\}$, the dropout in $\left\{0.2,0.5,0.8\right\}$. Furthermore, to keep consistency, the test tasks are identical for all baselines.
@@ 4.7 Ablation Study
In this section, we conducted an ablation study to verify the effectiveness of the motivation in CPT. Firstly, we remove the second stage from CPT and retained only the first stage, using the default meta-learning strategy to train until convergence. We refer to this variant as CPT w/o -SS. Secondly, we remove the first stage and keep the second stage, directly using the initial model to learn gradually difficult tasks. This variant is referred to as CPT w/o -FS. Finally, we reverse the order of the first and second stages in CPT. In this variant, we first learn hard tasks and then learn easy tasks, and we refer to it as CPT-reverse. The specific experimental results are shown in Table 3.
@@
Through the results of the ablation experiments, we have made the following findings:
@@ 5 Related Work
In this section, we briefly review some existing few-shot learning methods and curriculum learning methods.

### v4

@@ 1 Introduction
In this paper, we propose Competence-Progressive Training strategy for few-shot node classification, named as CPT. CPT aims to provide tasks of different difficulty levels for meta-learners at different stages to adapt to their current abilities and alleviate the issue of suboptimal points. Specifically, CPT uses the two-stage strategy Bengio et al. (2009); Hu et al. (2022) to train a meta-learner. In the first stage, CPT follows previous research by randomly sampling nodes to construct meta tasks and then employs the default training strategy(meta training), i.e., simply minimizing the given supervised loss, to produce a strong base meta-learner to start with. Although the base meta-learner may perform well on easy tasks and has gained some knowledge, it no longer benefits or improves its abilities from these tasks, and it might converge to suboptimal points. So the second stage focuses on generating more difficult tasks and fine-tuning it. Specifically, CPT increases the difficulty of each task by dropping the edges of its graph. The underlying motivation for this approach is to create a larger number of ’tail nodes’ within the meta-training tasks. This alteration is intended to enrich the meta-learning model with additional ’tail node’ meta-knowledge, as these nodes, with their fewer connections, typically present unique challenges in learning processes. Finally, CPT finetunes the base meta-learner by minimizing the loss over the increased hard task data. CPT can be easily implemented in the existing meta-learner training pipeline on graphs, as demonstrated in Algorithm 1. Additionally, CPT is compatible with any GNN-based meta-learner and can be employed with any desired supervised loss function.
@@ 2 PRELIMINARIES
2.1 Problem Definition
In this work, we focus on few-shot node classification (FSNC) tasks on a single graph. Formally, let $G=(\mathcal{V},\mathcal{E},X)=(A,X)$ denote an attributed graph, where $A=\mathop{\{0,1\}}^{V\times V}$ is the adjacency matrix representing the network structure, $X=[x_{1};x_{2};...;x_{V}]$ represents all the node features, $\mathcal{V}$ denotes the set of vertices $\{v_{1},v_{2},...,v_{V}\}$, and $\mathcal{E}$ denotes the set of edges $\{e_{1},e_{2},...,e_{E}\}$. Specifically, $A_{j,k}=1$ indicates that there is an edge between node $v_{j}$ and node $v_{k}$; otherwise, $A_{j,k}=0$. The few-shot node classification problem assumes the existence of a series of node classification tasks, $\mathcal{T}=\{T_{i}\}_{i=1}^{I}$, where $T_{i}$ denotes the given dataset of a task, and $I$ denotes the number of such tasks. FSNC tasks assume that these tasks are formed from target novel classes (i.e., $\mathcal{C}_{\text{novel}}$), where only a few labeled nodes are available per class. Additionally, there exists a disjoint set of base classes (i.e., $\mathcal{C}_{\text{base}}$, $\mathcal{C}_{\text{base}}\cap\mathcal{C}_{\text{novel}}=\emptyset$) on the graph, where the number of labeled nodes is sufficient. For better understanding, we first present the definition of an N-way K-shot node classification task as follows:
Definition 1. N-way K-shot Node Classification: Given an attributed graph $G=(A,X)$ with a specified node label space $\mathcal{C}$, $|\mathcal{C}|=N$. If for each class $c\in\mathcal{C}$, there are K labeled nodes (i.e., support set $\mathcal{S}$) as references and another R nodes (i.e., query set $\mathcal{Q}$) for prediction, then we term this task as an N-way K-shot Node Classification task.
Definition 2. Few-shot Node Classification: Given an attributed graph $G=(A,X)$ with a disjoint node label space $\mathcal{C}=\{\mathcal{C}_{\text{base}},\mathcal{C}_{\text{novel}}\}$. Substantial labeled nodes from $\mathcal{C}_{\text{base}}$ are available for sampling an arbitrary number of $N$-way $K$-shot Node Classification tasks for training. The goal is to perform $N$-way $K$-shot Node Classification for tasks sampled from $\mathcal{C}_{\text{novel}}$.
Few-shot Node Classification extends N-way K-shot Node Classification by utilizing labels from base classes for task training, while evaluation is performed using labels from novel classes. This approach assesses the model’s capacity to generalize and predict previously unseen categories.
2.2 Meta-Training on Graph
@@ 4.3 Implementation Details
Our methodology is formulated utilizing the PyTorch [^1] https://pytorch.org/ framework, and all subsequent evaluations are executed on a Linux-based server equipped with 1 NVIDIA Tesla V100 GPUs. For the specific implementation setting, we set the number of training epochs $T$ as 2000, the weight decay rate is set as 0.0005, the loss weight $\gamma$ is set as 1, the graph layers are set as 2 and the hidden size is set as in $\left\{16,32\right\}$. The READOUT function is implemented as mean-pooling. The other optimal hyper-parameters were determined via grid search on the validation set: the learning rate was searched in $\left\{0.03,0.01,0.005\right\}$, the dropout in $\left\{0.2,0.5,0.8\right\}$. Furthermore, to keep consistency, the test tasks are identical for all baselines.
@@ 4.6 Ablation Study
In this section, we conducted an ablation study to verify the effectiveness of the motivation in CPT. Firstly, we remove the second stage from CPT and retained only the first stage, using the default meta-learning strategy to train until convergence. We refer to this variant as CPT w/o -SS. Secondly, we remove the first stage and keep the second stage, directly using the initial model to learn gradually difficult tasks. This variant is referred to as CPT w/o -FS. Finally, we reverse the order of the first and second stages in CPT. In this variant, we first learn hard tasks and then learn easy tasks, and we refer to it as CPT-reverse. The specific experimental results are shown in Table 3. Through the results of the ablation experiments, we have made the following findings:
@@
4.7 Effectiveness in Mitigating Suboptimal Solutions.
@@
4.8 Effect of Meta-training Support Set Size
we are considering adding some experimental analyses, such as the impact of the size of the support set on performance. As the support set size $St$ increases, our method continues to be effective, as shown in Figure 5. We observe that under different $St$ settings, our method consistently outperforms all baselines.
Figure 5: Loss trend comparison of CPT, CPT w/o -SS, and CPT w/o -FS.